OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4.

OpenStudy (anonymous):

First you'll want to expand out the brackets.

OpenStudy (pawanyadav):

On solving 3x+3=-2x+2-4 5x=-2-3 X=-1.